OpenStudy (anonymous):

Sketch the graph of the function f defined by: f(x)=x if 0(less than or equal) X (less than or eqaul) 1 and f(x)=1 if x>1 Find integral from 0 to 2 f(x)dx. Why would the answer not be 2?

OpenStudy (zarkon):

\[\int\limits_{0}^{2}f(x)dx=\int\limits_{0}^{1}f(x)dx+\int\limits_{1}^{2}f(x)dx\]

OpenStudy (anonymous):

Thank you! Cant believe I didn't realize that!!